find a11 of an arithmetic sequence when a1=8 and the common difference d= -3
Answer #1

ai=8, Common differenced-3 all=? nth term of the asithmetic sequence : anz a+(n-1)d all= a + ( -1) all >8+ (10)(-3) =8-30